OpenStudy (anonymous):

Evaluate: -3 - (-6) Is it -9?

OpenStudy (anonymous):

its 3

rishavraj (rishavraj):

nope \[-ve \times -ve = +ve\]

OpenStudy (anonymous):

3-(-6) turns into: -3+6

OpenStudy (anonymous):

oh ok so i shoulded subtracted instead of multiplying thnx :)

OpenStudy (anonymous):

Its 3 trust me

OpenStudy (anonymous):

*shouldve

OpenStudy (anonymous):

two multiplied negatives = a positive bro

OpenStudy (anonymous):

my bad its positive 3 not -3

rishavraj (rishavraj):

- 3 - (- 6) = -3 + 6 = 3

OpenStudy (anonymous):

thnx guys :)

OpenStudy (anonymous):

See it's 3 like i said